Circular No. 5773

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993J IN NGC 3031 G. G. Pooley and D. A. Green, Mullard Radio Astronomy Observatory,

Cambridge, communicate: “Continued daily monitoring of SN 1993J with the Ryle Telescope at 15.5 GHz shows an approximately linear increase in flux density from its detection at this frequency on Apr. 5 (IAUC 5751) to ∼ 33 mJy during Apr. 21.70-21.85 UT. The source is clearly above the 20-mJy peak predicted for Apr. 19 on IAUC 5762, and does not yet appear to have reached a maximum. The position of the source is consistent with the radio positions reported on IAUC 5759, 5763, and 5768 and with the recent optical position on IAUC 5767.”

W. Romanishin, University of Oklahoma, reports observations of SN 1993J using the Kitt Peak 1.3-m telescope (+ SQIID infrared imaging detector). The following magnitudes (± 0.02 mag or less) have been derived relative to the brightest two comparison stars given by Odewahn (IAUC 5760): Apr. 14.17 UT, J = 10.22, H = 10.15, K = 9.91; 15.15, 10.18, 10.15, 9.87; 16.14, 10.13, 10.09, 9.89; 17.13, 10.10, 10.07, 9.83; 18.14, 10.05, 10.05, 9.82; 19.14, 10.04, 10.04, 9.82.

GRS 1915+105 I. F. Mirabel, Service d’Astrophysique, Centre d’Etudes de Saclay;

L. F. Rodriguez, Universidad Nacional Autónoma de México; J. Marti, Department of Astronomy and Meteorology, University of Barcelona; R. Teyssier and J. Paul, Service d’Astrophysique, Centre d’Etudes de Saclay; and M. Aurière, Observatoire du Pic du Midi, communicate: “Inside the 3′.5 SIGMA error radius (Cordier et al., Compton Symp., St. Louis) of this x-ray transient, discovered by WATCH onboard GRANAT (IAUC 5590) and observed with a slow rise to maximum intensity by BATSE on the Compton Observatory (IAUC 5619), we have detected with the Very Large Array a radio-counterpart candidate at α = 19^{h}12^{m}49^{s}.97, δ = +10^{o}51′25′′.6 ± 0′′.5 (equinox 1950.0). This source is time-variable: on 1992 Dec. 11, the 20-cm peak flux was 2.5 ± 0.1 mJy; on 1993 Mar. 10, it was 5.0 ± 0.1 mJy. R-band optical images obtained with the 2-m telescope at Pic du Midi during Apr. 19 and 20 show no star brighter than mag 21 at the position of the compact radio source. The nearest stars are > 4′′ away from the radio counterpart. Further radio observations at higher frequencies, as well as infrared observations in the H, J, and K bands, are in progress.”

1993 April 23 Daniel W. E. Green